A pizza shop has a staff of 25 workers who deliver 175 pizzas in a day. Today, 5 of your workers have fallen sick and have taken a holiday. What is the probability that you will be able to service an increased demand / delivery of 25% today through the remaining workers?


1
Expert's answer
2022-05-27T14:27:27-0400

Total no.of workers = 25

No.of pizzas delivered by 25 workers = 175

No.of pizzas delivered by each worker = 175/25 = 7


If 5 workers have fallen sick, then total no.of workers at present are 20.

Then total no.of pizzas delivered by 20 workers is "175(1+0.25) = 218.75"


No.of pizzas delivered by each worker is 218.75/20 = 10.9375

Then required probability is 10.9375/218.75=0.05